• Welcome to SpeedSolving.com — the world’s largest puzzle community!

    You’re currently browsing as a guest, which means you have limited access to discussions, resources, member profiles, and community features.

    Join 50,000+ cubers and puzzle enthusiasts from around the world to ask questions, share solves, improve faster, and be part of the community.

    Registration is free, fast, and easy — create your account today.

    Already a member? Log in here to hide this message and start participating.

Probability Thread

11-layered, not 13. We already have:
- megaminx: 3 layers
- gigaminx: 5 layers
- teraminx: 7 layers
-petaminx: 9 layers.
So examinx should have 11.

Well, you may think so, but an Examinx is not the next step up. By the regular orders of magnitude, it goes peta- exa- zetta-, but the one person who actually designed it named it wrong. Also, gigaminxes have 5 layers, not 6. I believe a 6 layered minx is called a super kilominx or star minx or something.

Examinx:
2WcYa.jpg

Based on Cmowla's formula, a super 13 layered minx would have 1.1*10^2934 combinations.
 
Last edited:
Well, you may think so, but an Examinx is not the next step up. By the regular orders of magnitude, it goes peta- exa- zetta-, but the one person who actually designed it named it wrong.
If the one person who designed it has no clue about how that naming scheme works, we're not forced to stick to his error forever just because he chose the wrong name in the first place. His puzzle is a zettaminx, not an examinx, period.

goodatthis said:
Also, gigaminxes have 5 layers, not 6.

Yes, U know, that was a typo, sorry.
 
What are the odds of getting a skip of orientation of the 8 edges other than the cross pieces? Because that's what just happened. I just got a solve that was rotationless, and was <RUL> except for one F move and one D move. This is a CFOP solve, not a ZZ solve. http://tinyurl.com/n5l3hy4

EDIT: actually, that's easy. The answer to getting a L8E orientation skip is just 1/2^7 or something.
EDIT2: ninja'd

What about the odds of having a scramble that's solvable with only two moves that aren't RUL? I guess another way to ask that would be what are the odds of having an EOline solvable in 2 non RUL moves?
 
Last edited:
What are the odds of getting a skip of orientation of the 8 edges other than the cross pieces? Because that's what just happened. I just got a solve that was rotationless, and was <RUL> except for one F move and one D move. This is a CFOP solve, not a ZZ solve. http://tinyurl.com/n5l3hy4

I believe 1/2^7 = 1/128 = 0.78%. (Each edge has two possible orientations, so 1/2 for each of 7 edges, and the 8th edge's orientation is defined by the other 7).
 
I was doing some skewb solves, when suddenly I had a last-5-centers skip. I solved 1 center + 4 corners first, then solved the last 4 corners with R' F R F', and then suddenly the skewb was solved. What is the probability for such a skip?
 
I was doing some skewb solves, when suddenly I had a last-5-centers skip. I solved 1 center + 4 corners first, then solved the last 4 corners with R' F R F', and then suddenly the skewb was solved. What is the probability for such a skip?
If you are left with 5 remaining centers, the U center can be in one of 5 places, so the chance of it being solved is 1/5. Given that the U center is solved, the F center can be in one of 4 places, so it has a 1/4 chance of being solved. Given the U and F centers are solved, the R center can be in one of 3 places, so it has a 1/3 chance of being solved. Once the U, F, and R centers are solved, L5C must be solved because there are two left and skewb can't have parity, so the chance of having a last five centers skip is 1/5*1/4*1/3 = 1/60. Another way to think of it is that there are 5 centers, so you could arrange them in 5! = 120 ways, but dividing by 2 (because skewb can't have parity) gives you 60 combinations, 1 of which is solved.

If you know Sarah's intermediate method, this is reduced to 1/12 because the U center is solved during the second step.
 
So this is a tricky one.

What are the chances that 3/4 of the 4 f2l slots are simple 3 move inserts+AUF in between? Lets just assume we can optimally solve pair order to force 3 move inserts. But no multislotting algs, simply different pair order.
 
If you had a set of 3x3 stickers (9 stickers of each sticker color, 6 colors), and you randomly stickered a cube, what are the odds that it will be solvable? ie, anytime there are 9 stickers of the same color on every face, it's 'solved', regardless of the color scheme.

What if we restricted the stickering such that centerpiece stickers must go on center pieces, corner stickers must go on corners, and edge stickers must go on corners (with a fitted sticker set or something in reality)?


What if we restricted 'solved' to only one specific color scheme?
 
Last edited:
Back
Top